OpenStudy (anonymous):

Find an equation of the line parallel to 6x + 2y = 49 containing the point (4,9).

OpenStudy (anonymous):

geometry or algebra1?

OpenStudy (anonymous):

Give me a point that is related to 4,9 such as...6,11 or something..

OpenStudy (anonymous):

adv algebra

OpenStudy (anonymous):

how about (6,110! lol @SexyBeast890

OpenStudy (anonymous):

My fault o.o It was 6,2 not 6,9 o.o Any who all we gotta do is change the =... So we keep: 6x + 2y But change: =49 = 56 so our equation will now look like 6x + 2y = 56

OpenStudy (anonymous):

So how would I write that in slope intercept form?

OpenStudy (anonymous):

since the lines are parallel, the slope will be the same. you also know one point on the line (4,9). then you plug these values into the line equation y=m(x-x1) + y1 where x1=4, y1=9 and m=slope=6

OpenStudy (anonymous):

^

OpenStudy (nikato):

6x+2y=49 first find the slope of this equation

OpenStudy (anonymous):

So what would the equation be in slope form?

OpenStudy (nikato):

point-slope or slope-intercept?

OpenStudy (anonymous):

slope-int

OpenStudy (anonymous):

subtract 6x from both sides then divide by 2y right?

OpenStudy (nikato):

ok. remember slope-intercept is y=mx+b where m is slope and b is your y-intercept

OpenStudy (anonymous):

that way you have it in y = mx + b

OpenStudy (nikato):

so 6x+2y=49 is in standard form, Ax +By=C and to find the slope, its -A/B so the slope is -6/2= -3

OpenStudy (nikato):

so you know m=-3 and (x,y)= (4,9) plug it in 9=-3(4)+b 9= -12+b +12 +12 ---------- 21=b so you're slope inter form is y=-3x+31

OpenStudy (nikato):

i meant y=-3x+21

OpenStudy (anonymous):

So 6x + 2y =49 as a parallel line containing the point (4,9) is y=-3x + 21?

OpenStudy (nikato):

yes

OpenStudy (nikato):

y=-3x+21 is the equation of the line that contains (4,9) and is parallel to 6x+2y=49

OpenStudy (anonymous):

THANKS! THIS MAKES SO MUCH MORE SENSE NOW!

OpenStudy (nikato):

you're welcome!